# Machine-assisted classification of potential biosignatures in earth-like exoplanets using low signal-to-noise ratio transmission spectra

### v1

@@ 5 Generation of training data
To train our algorithm, we need to generate a large sample of synthetic transmission spectra and their corresponding realizations with different levels of noise. For this purpose, we use the TauREx 3 framework (Al-Refaie et al., 2021). Given the challenge of creating a proper set of samples with different mixing ratios of fill and biosignature gases, atmospheric temperatures, pressure, and instrumental noise, we developed an independent Python package called MultiREx. The package is publicly available [^1] http://pypi.org/project/multirex. All the scripts created for our numerical experiments and to generate the plots in this paper are available in the GitHub public repository of the package [^2] https://github.com/D4san/MultiREx-public.
The synthetic transmission spectra were generated with the planetary properties of TRAPPIST-1 e, using the parameters provided by Agol et al. (2021) (see Table 3). The spectrum of TRAPPIST-1 was obtained from the Phoenix model grids (Husser et al., 2013). The atmospheric model implemented consists of 100 layers, with an isothermal temperature profile featuring three possible temperatures: 200 K, 287 K, and 400 K, with a common base pressure of $10^{5}$ Pa ($\approx 1$ atm) and a top pressure of $10^{-3}$ Pa.
Table 3: Parameters of planet and star (Agol et al., 2021) .
@@
Atmospheres were created with N2 as the fill-gas, and a fixed mixing ratio of CO2 of $10^{-2}$. This apparently arbitrary concentration is a compromise between the low levels observed on present Earth and a more primitive atmosphere (see e.g. Lin et al. 2021). It should also be noticed that the higher the level of CO2, the harder is the detection, using the traditional retrieval algorithms, of some interesting biosignatures such as ozone. However, and independently of the numerical experiments we describe in the following sections, we have ran tests using higher concentrations CO2 and noticed that, even above this level, the effects of the molecule on our results are negligible.
In addition to the fill-gases, we include, depending on the case, three additional molecules: CH4, O3, and H2O. These molecules were added such that an atmosphere could contain none, one, two, or all three of the molecules simultaneously. The mixing ratio of each molecule was selected among 10 possible values, all of them distributed log-uniformly between $10^{-10}$ and $10^{-1}$.
@@
In total, 3993 synthetic spectra were generated: [10$\times$10$\times$10 (all molecules included) + 3$\times$10$\times$10 (three combinations of 2 molecules) + 3$\times$10 (atmospheres with a single species) + 1 (atmosphere with only fill-gases)] $\times$ 3 (values of temperature).
@@
To train the algorithms, a constant Gaussian noise was added across the entire wavelength range depending on the SNR value. As a result, the 3993 synthetic spectra became many more (in some cases more than 100 times more). We trained the algorithms using spectra with six discrete SNR values: noise-less (for simplicity in notation, we will refer to this as 0), 1, 3, 6, 10, and 20. Additionally, for the training process, all spectra were normalized based on their maximum and minimum signal. For testing, however, we used a fine-grained interval of SNR.
In Figure 2, we show a representative transmission spectra, containing all interesting chemical species in concentrations similar to those of a Neoproterozoic Earth (Kaltenegger et al., 2020). We can see that many peaks can become obscured by other molecules if the mixing ratio and noise of the confounding molecule are increased.
To represent the strength of the noise in the simulated transmission spectrum in Figure 2, we include a bar over the strongest CO2 peak, showing the magnitude of the relative error. As expected for an SNR of 3, the noise can be as one third as the peak itself.
Figure 2: Absorption contributions from different molecules in a potentially inhabited TRAPPIST-1 e spectrum with surface concentrations similar to a Neoproterozoic Earth (Kaltenegger et al., 2020) and a mixing ratio of H2O fixed at $10^{-2}$. Each color represents a molecule (see legend). Error bars illustrate noise based on the SNR, using the CO2 strongest peak as the signal value.
@@ 6 Biosignature Random Forest Classification
Various classification experiments utilizing RF were conducted to investigate the potential for biosignature detection:
@@
    The first and simpler experiment was aimed at identifying interesting targets, namely planets which could actually have biosignatures and thus deserve further investigation. We call this experiment and the resulting model the Binary Classificator (BC).
@@
    The second experiment is more complex. It aims at classifying the planets by specific biosignatures, namely, determining if a planet could contain, for instance, ozone or not. We call this the Multilabel Classificator (MC).
@@
    The third experiment seeks to train what we call specialist algorithms. These algorithms are trained on planets having a specific biosignature molecule while they are used to classify the entire synthetic spectra dataset. We call this Specialised Classificator (SC).
For the implementation and performance evaluation of the RFs, the diverse functions offered by scikit-learn (Pedregosa et al., 2011) were employed. This approach allowed for a comprehensive analysis of biosignatures, leveraging the robust machine learning tools available within the mentioned library.
When applying RF classification, one crucial aspect is setting the threshold for classifying a planet as positive or negative. Although a conventional threshold of 0.5 is typically used, we adjusted it to increase the Recall due to the significant need to differentiate between FP and FN. This adjustment initially led to an increase in FP, prompting further lowering of the threshold. Ultimately, we optimized the threshold to ensure that the TNR remained above 0.6 for any SNR.
6.1 Binary Classificator
@@
This general experiment required the most diverse set of training and test spectra of all the experiments we performed. The distribution of planets having different combinations of biosignatures is summarised in Table 4. As we explained before, we generate synthetic spectra with 6 levels of SNR (rows in the table). The binary codes at the head of the table represent the presence (1) or absence (0) of CH4, O3, and H2O, respectively. For instance, the third column, labelled 010, corresponds to planets without CH4 (0), containing O3 (1) and in the absence of H2O (0). The number in each cell indicates how many times a noisy spectrum was generated for that type of atmosphere.
Table 4: Training data distribution for the RF classificator focused on interesting classification. A SNR of 0 correspond to a noise-less spectra.
SNR 000 100 010 001 110 101 011 111
1 20000 500 300 1000 20 20 20 5
3 20000 500 300 1000 20 20 20 5
6 20000 500 300 1000 20 20 20 5
10 10000 500 300 500 20 20 20 5
20 10000 500 300 500 20 20 20 5
0 10000 500 300 500 20 20 20 5
@@
In the example before, for training the algorithm using synthetic spectra with 010 composition and having SNR of 10 (fourth row) we generate 300 noisy spectra per each value of O3 concentration (10 in total) and per each atmospheric temperature (3 in total). This gave us a total of 9000 spectra for this composition alone. Similarly, for the case of planets having all molecules (111), we generate 5 noisy spectra per each molecule concentration ($10\times 10\times 10$), and per temperature value (3), ending up with 15000 spectra for this composition.
A total of 195 000 synthetic spectra was used in this experiment.
@@
It is important to stress that the quantity and disparity of data labelled as bio or non-bio can significantly influence RF performance, especially for low SNR signals. Adjusting the training data proportion towards more negative or positive examples as necessary can enhance the model’s ability to recognise the presence or absence of certain molecules under challenging noisy conditions.
@@
Using this data, we trained a RF classificator to determine whether new spectra are interesting or not-interesting. In this context, an interesting spectrum is defined by the potential presence of CH4 or O3, while the presence of H2O is treated as interference rather than a biosignature. The training was conducted using 80% of the data, and an initial validation was performed with the remaining 20%.
@@
Our RF model contained 400 estimators, used an entropy criterion, and a minimum samples leaf of 3. After training, the classification thresholds were set at 0.4.
@@
The resulting trained RF was tested using a set of 180 000 new synthetic spectra per SNR value, equally distributed between bio and non-bio categories. The non-bio spectra consisted of 45 000 planets with CO2 and N2, and 45 000 with H2O, CO2 and N2. The bio spectra had 15 000 samples for each possible combination of biosignature molecules (6 combinations) which gives 90 000 total test spectra.
@@ 6.1.1 BC metrics
In Figure 3, we show the value of each metric defined in subsection 2.2 as a function of SNR for the trained algorithm. It is important to stress that these results are obtained when classifying spectra of all possible mixing ratios. Therefore, this metric will provide a conservative measurement of the algorithm performance. If we would focus only on the planets with the highest concentrations of the biosignatures gasses, the results will be much better.
As expected, as SNR increases, most metrics improve. Even in the worst-case scenario of SNR=1, all metrics have a value above 0.6, which means that even with the noisiest spectra, it works better than a blind guess.
Figure 3: Global metrics of binary classification of interesting (with biosignatures) and non-interesting planets across different SNR values. The shaded area below 0.6 in the graphs indicates suboptimal classification performance.
Interestingly, the algorithm can classify non-bio as non-interesting planets very efficiently (TNR $\approx 1$) even for values of SNR as low as 3. This means that the algorithm could be very good at avoiding wasting time on uninteresting planets. This is confirmed by the behaviour of the Precision metric.
An interesting behaviour is noticed in the Recall metric, where the algorithm seems to trade-off the improvement/worsening in TNR at the expense of Recall. In other words, for low values of SNR, the algorithm becomes more pessimistic (discarding bio planets as not interesting) as SNR increases, while at the same time, it avoids classifying as interesting planets that are not (high TNR).
In summary, the analysis of the metrics shows that our BC algorithm performs relatively well at labeling as interesting planets that have biosignatures, but it performs much better for discarding as non interesting planets without biosignatures. This will avoid wasting valuable research time on non interesting targets.
@@ 6.1.2 BC minimum mixing ratio for detection
As a secondary analysis, for each value of the SNR, we sorted the classified synthetic spectra according to the CH4 mixing ratio. Then, we selected the spectra having a specific CH4 mixing ratio and calculate our binary classification metrics only for planets with this concentration.
As expected, for low values of the mixing ratio, the Recall metric was generally low. We repeated this procedure until the Recall reached a value of 0.6. The corresponding mixing ratio is called the minimum mixing ratio for detection. We repeated this analysis for O3.
In Figure 4, we show the minimum mixing ratio for the detection of each biosignature as a function of SNR values. In the case of CH4, it is observed that the algorithm, with an SNR of 1, successfully classifies as interesting atmospheres with a mixing ratio as low as 10-6. For comparison, the present Earth has a mixing ratio of the order of $10^{-6}$, but in the past it has reached levels as high as $10^{-3}$ or larger (Kaltenegger et al., 2007, 2020).
The capability of the BC algorithm for detecting low levels of CO2 is lost when SNR $>2$ and only recovers when SNR $>5$. This trend is related to the decrease in Recall within this SNR range, as shown in Figure 3.
@@
The case of O3 is more complex. Only when the SNR reaches values larger than 10-15 (that in the case of TRAPPIST-1 e would require with JWST more than 100 transits) the minimum surface mixing ratio detection for this biosignature is of the order of that on present Earth. However, if a planet has a global biosphere with levels of CH4 comparable to that of Earth, even if the levels of O3 are low, our algorithm will tag the planet as interesting. In other experiments we will show that the detection capabilities for ozone can be improved by training algorithms for multilabel classification (see subsection 6.3).

### v4

@@ 6 Generation of training data
To train our algorithm, we need to generate a large sample of synthetic transmission spectra and their corresponding realizations with different levels of noise and stellar contamination. For this purpose, for generating the transmission spectra we use the TauREx 3 framework (Al-Refaie et al., 2021) and, stellar contamination effects were simulated using the POSEIDON package (MacDonald & Madhusudhan, 2017; MacDonald, 2023).
Due to the complexity involved in generating a comprehensive dataset with varying mixing ratios of fill and potential biosignature gases, atmospheric temperatures, pressures, and instrumental noise, we developed an independent Python package called MultiREx. The package is publicly available [^1] http://pypi.org/project/multirex. All scripts created for our numerical experiments and to generate the plots in this paper are available in the GitHub public repository for the package [^2] https://github.com/D4san/MultiREx-public.
The synthetic transmission spectra were generated using the planetary properties of TRAPPIST-1 e, with the parameters provided by Agol et al. (2021) (see Table 3). The spectrum of TRAPPIST-1 was obtained from the Phoenix model grids (Husser et al., 2013). The atmospheric model implemented consists of 100 layers, with an isothermal temperature profile featuring three possible temperatures: 200 K, 287 K, and 400 K, with a common base pressure of $10^{5}$ Pa ($\approx 1$ atm) and a top pressure of $10^{-3}$ Pa.
Table 3: Planetary and stellar parameters as provided by (Agol et al., 2021).
@@
Although we are training our algorithms to detect planets with potential biosignatures, including atmospheres that are not necessarily conducive to life—such as those with freezing temperatures of 200 K or infernal conditions at 400 K, which are incompatible with our target model—the training of a machine-learning algorithm requires the most diverse set of conditions achievable in our models. Indeed, training the algorithms with a narrow range of temperatures, either strictly compatible with life or matching the atmospheric composition of the model planet, may lead to overfitting.
Atmospheres were created with N2 as the fill gas and a fixed mixing ratio of CO2 of $10^{-2}$. This concentration was selected as an intermediate scenario between the low CO2 levels observed in Earth’s current atmosphere and those predicted for more primitive atmospheric conditions (see, e.g., Lin et al. 2021). It should also be noted that the higher the level of CO2, the harder it becomes to detect, using traditional retrieval algorithms, certain interesting biosignatures/bioindicators such as ozone. However, independent of the numerical experiments described in the following sections, we have run tests with higher concentrations of CO2 and observed that, even above this level, the molecule’s effects on our results are negligible.
In addition to the fill gas, we include, depending on the case, three additional molecules: CH4, O3, and H2O. These molecules were added such that an atmosphere could contain none, one, two, or all three simultaneously. The mixing ratio of each molecule was selected from 8 possible values, all of which are distributed log-uniformly between $10^{-8}$ and $10^{-1}$.
Additionally, we included in all the training sets spectra of airless planets (i.e. planets with a flat transmission spectrum). These control scenarios serve as baselines to assess the performance of classification algorithms designed to detect molecular species where no atmospheric features are present. To account for stellar contamination, we calculated the $\epsilon$ function (see Equation 7) using POSEIDON (MacDonald & Madhusudhan, 2017; MacDonald, 2023). For this procedure, we follow the method outlined by Rackham et al. (2018), who derived spot and facula temperature parameterisations as well as coverage fractions for the TRAPPIST-1 system.
For the spot and faculae temperatures, we used fixed values of $T_{\text{spot}}=0.86\,T_{\text{phot}}$ and $T_{\text{fac}}=T_{\text{phot}}+100\,\mathrm{K}$, where $T_{\text{phot}}$ represents the stellar photospheric temperature. For the coverage fractions, we based our samples on the values reported by Rackham et al. (2018): $f_{\text{spot}}=8^{+18}_{-7}\%$ and $f_{\text{fac}}=54^{+16}_{-46}\%$. Nine different combinations of $f_{\text{spot}}$ and $f_{\text{fac}}$ (minimum, mean, and maximum values for each), plus one uncontaminated scenario, were used when contaminating the synthetic transmission spectra. As a result, for each planetary spectrum (i.e. atmospheric temperature, composition, mixing ratio), we generate a total of ten possible contaminated spectra.
@@
In total, 2188 synthetic spectra were generated: [8$\times$8$\times$8 (all molecules included) + 3$\times$8$\times$8 (three combinations of two molecules) + 3$\times$8 (atmospheres with a single species) + 1 (atmosphere with only fill gases)] $\times$ 3 (temperature values) + 1 (flat spectrum). By applying the ten stellar-contamination scenarios (nine contaminated + one uncontaminated) to each of these base spectra, the overall dataset was expanded to $21,880$ distinct synthetic spectra.
@@
To train the algorithms, we added to the theoretical spectra a constant Gaussian noise across the entire wavelength range. The noise was calculated according to the assumed SNR value. As a result, the number of synthetic spectra is substantially increased with respect to the 21,880 theoretical set. In some cases by a factor of 100 or more.
We trained the algorithms using five discrete SNR values: noise-less (denoted 0), 1, 3, 6, and 10. Larger SNR values were not used since from our PandExo analysis, such a noisy spectra would require more than 100 transits with NIRSpec PRISM (see section 8.
For each set of atmospheric parameters, we determined the noise standard deviation from the signal of the uncontaminated spectrum, thus ensuring consistency among the contaminated versions of that family. Finally, all spectra were normalised between their minimum and maximum values for training. For testing, however, we used a fine-grained interval of SNR.
In Figure 3, we show a representative transmission spectrum containing all interesting chemical species in concentrations similar to those of a Neoproterozoic Earth (Kaltenegger et al., 2020). We can see that many peaks may become obscured by other molecules if the mixing ratio and noise level of the confounding molecule are increased.
To illustrate the strength of the noise in the simulated transmission spectrum in Figure 3, we include a bar over the strongest CO2 peak, indicating the magnitude of the relative error. As expected for an SNR of 3, the noise can be as high as one third of the peak amplitude.
Figure 3: Absorption contributions from different molecules in a potentially inhabited TRAPPIST-1 e spectrum with surface concentrations similar to a Neoproterozoic Earth (Kaltenegger et al., 2020) and a mixing ratio of H2O fixed at $10^{-2}$. Each colour represents a molecule (see legend). Error bars illustrate noise based on the SNR, with the strongest CO2 peak used to define the signal. The error bars represent a $1\sigma$ uncertainty.
For illustrating the effect of the level of noise in the spectra we deal with in this work, we show in Figure 4 several examples of signals at SNR=1. These figures illustrate well how featureless noisy spectra seem even at large mixing ratios of the interesting molecules.
Figure 4: Examples of synthetic transmission spectra for a TRAPPIST-1 e planet under noisy conditions at $\text{SNR}=1$. Each panel corresponds to a distinct combination of methane (CH4) and ozone (O3) mixing ratios, illustrating how noise at this level affects the observable spectral features. The top three panels (yellow, red, and blue points) show the noisy flux measurements across the wavelength range for three different CH4–O3 scenarios. The bottom panel compares these scenarios in a single line plot, highlighting the variation in spectral peaks and the extent to which noise can obscure molecular signatures.
@@ 7 Biosignature Random Forest Classification
Various classification experiments employing Random Forest (RF) were performed to investigate the potential for biosignature detection:
@@
    The first and simpler experiment aimed to identify “interesting” targets, i.e. planets that could potentially have biosignatures/bioindicators and thus merit further investigation. We refer to this experiment and the resulting model as the Binary Classifier (BC).
@@
    The second experiment is more complex. It seeks to classify planets by specific molecules (e.g. determining whether a planet could contain ozone). We call this the Multilabel Classifier (MC).
@@
    The third experiment involves training what we call “specialist” algorithms. These algorithms are trained on planets that have a specific biosignature molecule, yet they are applied to classify the entire synthetic spectra dataset. We refer to this as the Specialised Classifier (SC).
For the implementation and performance evaluation of the RFs, we used the diverse functionalities provided by scikit-learn (Pedregosa et al., 2011). This approach enabled a comprehensive analysis of biosignatures/bioindicators, leveraging the robust machine learning tools available within that library.
A critical step in RF classification is determining the optimal threshold for classifying a planet as positive or negative. Although a conventional threshold of 0.5 is typically used, we adjusted it to increase the Recall, given the significant need to distinguish between FP and FN. This adjustment initially led to an increase in FP, prompting further lowering of the threshold. The threshold was optimised to maintain the TNR above 0.6 across all SNR levels.
7.1 Binary Classifier
@@
This general experiment required the broadest range of atmospheric compositions among all performed experiments The distribution of planets having different combinations of biosignatures/bioindicators is summarised in Table 4. As we explained before, we generate synthetic spectra with 6 levels of SNR (rows in the table). The binary codes at the head of the table indicate the presence (1) or absence (0) of CH4, O3, and H2O, respectively. For instance, the third column, labelled 010, corresponds to planets without CH4 (0), containing O3 (1), and in the absence of H2O (0). The values in each cell represent the total number of noisy spectra generated for each atmospheric type.
Table 4: Distribution of training data for the RF classifier targeting interesting planets. Each row represents a different SNR level (with 0 indicating a noise-free spectrum), while each column is labelled with a three-digit binary code indicating the presence (1) or absence (0) of CH4, O3, and H2O, respectively (e.g., “010” means no CH4, presence of O3, and no H2O). The values in each cell (in thousands) denote the total number of noisy spectra generated for that specific atmospheric configuration.
SNR 000 100 010 001 110 101 011 111 airless
1 300 60 60 60 57.6 57.6 57.6 76.8 100
3 300 60 60 60 57.6 57.6 57.6 76.8 100
6 300 24 24 24 38.4 38.4 38.4 46.08 100
10 300 24 24 24 38.4 38.4 38.4 46.08 100
0 300 60 60 60 96 96 96 57.6 100
@@
In the example above, to train the algorithm using synthetic spectra with 010 compositions and an SNR of 10 (fourth row), we generated 250 noisy spectra for each O3 concentration value (8 in total) and for each atmospheric temperature (3 in total). Considering the 10 different contamination scenarios (9 contaminated + 1 uncontaminated), this resulted in 2500 spectra per O3 concentration and temperature combination, leading to a total of 60,000 spectra for this composition alone. Similarly, for the case of planets having all molecules (111), we generated 5 noisy spectra per molecule concentration (8$\times$8$\times$8), per temperature value (3), and per contamination scenario (10 in total), resulting in 76,800 spectra for this composition.
A total of 3,770,560 synthetic spectra were used in this experiment.
To avoid overfitting on the highest SNR spectra, we provided the algorithm with a smaller number of samples having SNR values of 6 and 10.
@@
It should be emphasised that the quantity and distribution of data labelled as bio or non-bio can significantly influence RF performance, especially for low SNR signals. Adjusting the training data proportion towards more negative or positive examples, as necessary, can enhance the model’s ability to recognise the presence or absence of certain molecules under challenging noisy conditions.
@@
Using this data, we trained an RF classifier to determine whether new spectra are interesting or not-interesting. In this context, an interesting spectrum is defined by the potential presence of CH4 or O3, while the presence of H2O is treated as interference rather than as a biosignature. The training was conducted using 80% of the data, with an initial validation performed on the remaining 20%.
@@
Our RF model contained 100 estimators, used the entropy criterion, and a minimum samples leaf of 3. After training, the classification threshold was set at 0.45.
@@
The resulting trained RF was tested using a set of approximately 900,000 new synthetic spectra per SNR value, equally distributed between the bio and non-bio categories. The non-bio spectra consisted of 150,000 planets with CO2 and N2, 150,000 planets without an atmosphere, and approximately 150,000 with H2O, CO2, and N2. The bio spectra comprised roughly 75,000 samples for each possible combination of biosignature/bioindicator molecules (6 combinations), giving a total of approximately 450,000 test spectra. As can be seen, the test set is balanced between samples with and without biosignature/bioindicator molecules.
@@ 7.1.1 BC metrics
In Figure 5, we show the value of each metric defined in subsection 2.2 as a function of SNR for the trained algorithm. It is important to stress that these results are obtained when classifying spectra with all possible mixing ratios. Therefore, this metric provides a conservative measure of the algorithm’s performance. If we were to focus only on the planets with the highest concentrations of the biosignature/bioindicator gases, the results would be much better.
As expected, as SNR increases, most metrics improve. Even in the worst-case scenario of SNR = 1, all metrics have a value above 0.6, indicating that even with the noisiest and most contaminated spectra (as represented by the shaded area), the algorithm performs better than a blind guess.
Figure 5: Global metrics (TNR, Recall, Precision, and F1 score; see Table 2 for definitions) for the binary classification of interesting (with biosignatures/bioindicators) and non-interesting planets across different SNR values. The pink shaded area below 0.6 in the graphs indicates suboptimal classification performance. The solid lines represent the results for spectra without stellar contamination, while the coloured shaded areas around each curve indicate the variance in the results produced by stellar contamination.
Notably, the algorithm can classify non-bio as non-interesting planets very efficiently (TNR $\approx 0.9$) even for values of SNR as low as 3. This means that the algorithm could be very good at avoiding wasting time on uninteresting planets. This is confirmed by the behaviour of the Precision metric.
The analysis of these metrics indicates that our BC algorithm performs adequately in labelling planets with biosignatures/bioindicators as interesting, but it performs even better at discarding planets without these features as non-interesting. This capability is important for preventing the unnecessary allocation of research time to targets that are unlikely to yield valuable biosignature information.
@@ 7.1.2 BC minimum mixing ratio for detection
As a secondary analysis, for each SNR value we sorted the classified synthetic spectra according to the CH4 mixing ratio. We then selected the spectra corresponding to a specific CH4 mixing ratio and computed our binary classification metrics solely for planets with that concentration. It is important to note that while Figure 5 presents the overall performance, it does not reveal how the classifier behaves at lower mixing ratios, as scenarios with high mixing ratios can mask the performance at the lower end.
For low mixing ratios, the Recall metric was generally low, consistent with expectations. We repeated this procedure until the Recall reached a value of 0.6; the corresponding mixing ratio is termed the minimum mixing ratio for detection. We performed this analysis for O3 as well, thereby providing a clearer insight into the classifier’s sensitivity at lower mixing ratio values.
In Figure 6, we show the minimum mixing ratio required for the detection of each biosignature as a function of SNR. In the case of CH4, it is observed that the algorithm, at SNRs of 1–2, successfully classifies as interesting atmospheres with mixing ratios as low as $10^{-6}$ to $10^{-4}$. For comparison, the present Earth has a CH4 mixing ratio on the order of $10^{-6}$, but in the past it has reached levels as high as $10^{-3}$ or greater (Kaltenegger et al., 2007, 2020). Notably, stellar contamination (indicated by the shaded areas in the figure) tends to lower the minimum mixing ratio required for detection. In other words, when a planetary transmission spectrum is heavily contaminated, our BC algorithm is more inclined to flag it as interesting. This effect may arise because contamination—especially from spots—can enhance certain spectral features, thereby aiding the algorithm in identifying the relevant molecules. Consequently, the presence of spots might also influence the autoencoder’s ability to reconstruct an uncontaminated spectrum, ultimately improving the BC algorithm’s results.
@@
The case of O3 is particularly interesting. Our BC algorithm can identify as interesting most planets with ozone levels on the order of those observed on present-day Earth ($10^{-5}$$10^{-6}$), starting at an SNR of 2. In the case of TRAPPIST-1 e, achieving such low SNR values would require only about 6 transits with JWST. It is important to note that although present-day ozone levels at Earth’s surface are around $10^{-8}$, transmission spectra sample an air column that includes layers with much higher ozone concentrations, effectively raising the observed value to $10^{-5}$$10^{-6}$ (see Table 2 in Lustig-Yaeger et al. 2023 and Figure 5 in Lin et al. 2021).
Finally, we observe that even when the levels of O3 or CH4 are relatively low in inhabited planets—i.e. planets exhibiting a non-equilibrium abundance of both molecular species—our algorithm still flags such planets as interesting, even if the detectability of one of the molecules falls below the minimum mixing ratio threshold. Conversely, if a planet exhibits only one of these molecules (and thus should be considered as having a bioindicator rather than a full biosignature), the BC algorithm will still label it as interesting. This observation motivates the development of multilabel algorithms capable of labelling the atmosphere according to individual species (see subsubsection 7.2.1 and subsubsection 7.3.1).